# Lazy-DaSH: Lazy Approach for Hypergraph-based Multi-robot Task and Motion Planning

### v1

@@ V-C2 Wall
Lazy-DaSH makes assumptions on motion feasibility that may turn out to be incorrect. We evaluate the cost incurred by the method in recognizing and adapting to invalid motion assumptions, by introducing a set of thin walls to sorting problem, which violates the assumptions made in the lazy stages of the method. In Wall, illustrated in Figs. 3(c) and 3(d), five manipulators are tasked with transferring the objects to designated boxes across the walls. Lazy-DaSH’s motion validation adopts an optimistic approach, which can result in infeasible motions being incorporated into the plan. To address this, the constraint feedback mechanism detects such infeasibility and triggers either a revised motion plan or necessary adjustments at the task planning level. To assess this capability, we progressively increase the number of walls, adding complexity to the problem. Manipulators near the wall are not only responsible for sorting the objects but also serve as intermediaries, assisting in passing the objects over the walls. The result shows that Lazy-DaSH scales more than twice as effectively as DaSH in terms of the number of objects in a presence of interaction obstructions.
@@ V-C3 Shelf-Wall
To demonstrate its ability to handle geometric constraints, DaSH [2] introduces the “Shelf” scenario, where objects must be placed on a shelf such that each rear object is completely blocked by the front object. We extend this concept with a more complex scenario, Shelf-Wall, which integrates the Shelf and Wall environments, treating each shelf as a barrier that obstructs both manipulator interactions and object access. As illustrated in Figs. 3(e) and 3(f), two or four robots are tasked with rearrange objects while satisfying the geometric constraints. For example, green (yellow) blocks must be positioned behind the red (blue) blocks and placed on the shelf facing the opposite side. Manipulators must hand over objects to robots on the opposite side, as the shelves act as a physical wall between them. As the number of robots increases, each object transfer requires at least two handovers, similar to the Sort scenario.
Since these geometric constraints are not accounted for in the initial optimistic schedule, the conflict resolution layer detects them and ensures that subsequent queries generate a valid schedule that respects these constraints. Specifically, the conflict resolution layer identifies hyperarc-vertex conflicts, such as when a manipulator’s motion (hyperarc) attempts to pass through a statically placed object (vertex)—for example, a green (yellow) block intersecting a fixed red (blue) block. The results demonstrate that Lazy-DaSH successfully handles 4 robots with 20 objects, whereas DaSH fails to find a solution within the time limit even for a 2 robots scenario.
V-C4 Lab
@@
To address problems where each object requires multiple operations to reach its goal state, we introduce the Lab scenario, inspired by the Mind in Vitro (MiV) project [29]. In this scenario, two 3-axis gantry robots inspect specimens within a sealed container, as illustrated in Fig. 3(g). The process begins with the robots removing the lid of the container and placing it in a predefined sanitization zone for sterilization. They then use tools attached to their end effectors to inspect the specimen. Once the inspection is complete, the lid is returned to the container to preserve the specimen’s environment. Due to the limited space in the sanitization zone, the planner must determine a valid placement that prevents lid collisions. This placement is constrained by vertex-vertex conflicts, where each vertex represents a potential lid position in the sanitization zone. The results demonstrate that Lazy-DaSH outperforms DaSH, successfully handling more than twice the number of objects requiring multi-stage operations.
@@ V-D Analysis and Discussion
This section analyzes and discusses the experiment results, focusing on four distinct capabilities of Lazy-DaSH using the criteria defined in Section V-A.
@@ V-D1 Overview
Across all tested scenarios, Lazy-DaSH demonstrated significantly improved scalability comared to DaSH, as illustrated in Figures 5 and 6. The scalability analysis in [2] shows that the number of objects leads to linear growth in representation size, while the number of robots results in quadratic growth. This makes DaSH particularly inefficient in environments with many robots, where the exhaustive expansion of the search space imposes substantial computational overhead. This limitation was evident in the Sort scenarios, which are designed to test scalability with respect to the problem size in terms of both the number of robots and objects.
DaSH also struggles in scenarios that require frequent replanning at both the task and motion levels. In environments with complex constraints, such as Wall, Shelf-Wall, and Lab scenarios, DaSH fails to scale because it focused solely on incorporating motion-level constraints without adapting the higher-level task structure. In particular, the Shelf-Wall scenario, which is heavily constrained by geometric constraints, was not solvable by DaSH even in its smallest environment setup. In contrast, Lazy-DaSH employs a hierarchical query framework that combines lazy motion validation with a constraint feedback mechanism. This design enables focused motion refinements, targeted backtracking, and high-level task replanning, thereby improving efficiency and scalability in complex environments.
@@ V-D2 Representation Size
The planning performance of both DaSH and Lazy-DaSH is directly influenced by the size of their representations, which serve as the foundation for iterative queries. We compare the size of $\mathcal{H}_{\mathcal{M}}$ in terms of the number of vertices and hyperarcs, as shown in Figure 5. The comparison of $\mathcal{H}_{\mathcal{T}}$ is omitted since Lazy-DaSH simply extends DaSH’s $\mathcal{H}_{\mathcal{T}}$ by incorporating additional object-only task space elements. The results show that the number of vertices and hyperarcs in $\mathcal{H}_{\mathcal{M}}$ is significantly smaller in Lazy-DaSH across all scenarios. Notably, the number of hyperarcs—critical due to their impact on motion validation checks—is up to two orders of magnitude smaller in Sort, Wall, and Lab, and up to one order of magnitude smaller in Shelf-Wall compared to DaSH. This reduction in representation size demonstrates the effectiveness of Lazy-DaSH’s task query phase in identifying only the necessary task space elements. Additionally, it highlights the efficacy of the constraint management system in selectively propagating only necessary constraints, ensuring a more compact and efficient representation. The impact of this reduction is further analyzed in the following discussion on overall planning performance.
@@ V-D3 Planning Time
Fig. 6 presents the total planning time along with a detailed breakdown. The first three columns represent the key components contributing to the total planning time: motion hypergraph construction, task and motion query, and conflict resolution. The fourth column shows the total planning time, which is the sum of these three components. To facilitate direct comparisons across different planning times, the y-axis for each scenario are kept consistent across all columns.
@@
The representation construction time includes both task and motion representation construction, with motion representation accounting for the largest portion due to the computational cost of collision checking. In DaSH, this process involves constructing a fully validated roadmap upfront. In contrast, Lazy-DaSH adopts a lazy construction approach, initially generating an edge-invalidated roadmap and incrementally refining it in response to failures encountered during motion queries or conflict resolution. In addition to the reduced representation size in Lazy-DaSH, this incremental approach significantly decreases construction time, achieving up to two orders of magnitude reduction in scenarios where both methods are comparable.
@@
The query times for DaSH and Lazy-DaSH differ in definition. In DaSH, it represents the combined task and motion planning query time, whereas in Lazy-DaSH, it is the sum of separate task and motion queries. In both methods, queries may be iteratively triggered by failures in planning or conflict resolution, and the cumulative time for these iterations is reflected in the total query time. A key observation is that as the number of objects increases, the query time in DaSH grows at a much steeper rate compared to Lazy-DaSH. This is due to DaSH’s representation size becoming intractable, leading to an exponential increase in search space. In contrast, Lazy-DaSH maintains a more scalable query process, significantly reducing query time. The reduction reaches up to three orders of magnitude in scenarios where DaSH is able to find a solution.
@@
Conflict resolution time, which involves the iterative detection and resolution of conflicts over entire motion paths, increasingly contributes to the total planning time as the number of objects grows. This is inherent to the conflict-based search approach [28], where motions are recomputed and entire paths are revalidated for collisions as conflicts are detected. As the length of the motion plan increases, typically due to an increased number of objects, the likelihood of recomputation and revalidation increases, becoming a primary source of computation in the overall planning process. Although this suggests that the conflict resolution layer could become a bottleneck in the planning process as the problem size increases, it is worth noting that this layer can be replaced with off the shelf schedulers or their variants to improve scalability. Future work will explore integrating alternative motion scheduling techniques into the conflict resolution layer to further enhance scalability.
@@ VI Conclusion
In this paper, we introduce Lazy-Decomposable State Space Hypergraph (Lazy-DaSH), a novel hypergraph-based approach to multi-robot object rearrangement that extends DaSH. Lazy-DaSH separates task and motion planning into distinct layers connected through a constraint feedback mechanism and employs a lazy motion evaluation strategy, validating only the motions relevant to the candidate task plan to minimize unnecessary computation. The constraint feedback mechanism manages infeasible motions resulting from lazy evaluation and incorporates task/motion constraints detected by the conflict resolution layer, enabling dynamic updates and refinements to the planning representation.
This approach allows the planner to construct a minimal yet sufficient representation while efficiently managing constraints critical for task completion. Experimental results across four scenarios demonstrate that Lazy-DaSH significantly improves scalability and planning speed in highly constrained environments compared to DaSH. Additionally, one of the experiments is validated through hardware implementation:https://youtu.be/bWQwy0dWAJI. Code is available at:https://github.com/parasollab/open-ppl/tree/lazy-hypergraph
Acknowledgement
This work was supported in part by the U.S. National Science Foundation’s “Expeditions: Mind in Vitro: Computing with Living Neurons” under award No. IIS-2123781, and by the IBM-Illinois Discovery Accelerator Institute and the Center for Networked Intelligent Components and Environments (C-NICE) at the University of Illinois.
Morales is supported in part by Asociación Mexicana de Cultura A.C.

### v2

@@ V-C2 Wall
This scenario is designed to evaluate how the planner adapts when its optimistic motion assumptions fail due to infeasible robot interactions. To create such conditions, we introduce thin walls into the Sorting setup, adding obstacles that invalidate the initial lazy assumptions. As illustrated in Figs. 5(c) and 5(d), eight manipulators must transport objects to designated boxes across these walls. To systematically increase complexity, we progressively add more walls. Robots positioned near a wall not only sort their own objects but also act as intermediaries, passing items across the obstructions to teammates on the other side.
The result shows that Lazy-DaSH scales up to 8 robots with more than 40 objects, while DaSH scales with up to 5 robots with 10 objects in a presence of interaction obstructions.
@@ V-C3 Shelf-wall
To demonstrate its ability to handle geometric constraints, DaSH [2] introduces the “Shelf” scenario, where objects must be placed on a shelf such that each rear object is completely blocked by the front object. We extend this concept with a more complex scenario, Shelf-wall, which integrates the Shelf and Wall environments, treating each shelf as a barrier that obstructs both manipulator interactions and object access. As illustrated in Figs. 5(e) and 5(f), two or four robots are tasked with rearrange objects while satisfying the geometric constraints. For example, green (yellow) blocks must be positioned behind the red (blue) blocks and placed on the shelf facing the opposite side. Manipulators must hand over objects to robots on the opposite side, as the shelves act as a physical wall between them. As the number of robots increases, each object transfer requires at least two handovers, similar to the Sorting scenario.
The results demonstrate that Lazy-DaSH successfully handles 4 robots with 32 objects, whereas DaSH fails to find a solution within the time limit even for a 2 robots scenario.
V-C4 Stocking
To evaluate the task space expansion and non-monotonicity capabilities of Lazy-DaSH, we consider a real-world stocking scenario in which two robotic arms place new items on a shelf behind existing ones (Fig. 5(g)). Since the existing items must remain in place, the task space must expand; otherwise, placing new objects behind them would be impossible. Moreover, the new objects must be taken out and restocked in a specific order to prevent obstruction. The new items to be stocked are initially placed on the far side of the shelf, similar to the Sorting and Shelf-wall scenarios, so robot handovers are still required.
We increase complexity by layering the stocked objects, thereby introducing different levels of non-monotonicity and geometric constraints. For example, in a two-layer configuration, the task space must expand for the front objects, which need to be temporarily moved out to place new items at the back. In a three-layer configuration, the task space expands for the objects in the first two rows, which must be moved out in sequence and then restocked in reverse order.
The experiments demonstrate that Lazy-DaSH successfully handles three layers of non-monotonicity with 2 robots and 18 objects, and two layers with 2 robots and 28 objects, validating both its scalability and its ability to resolve geometric constraints and non-monotonicity, where DaSH fails.
V-C5 Lab
@@
To address problems where each object requires multiple operations to reach its goal state, we introduce the Lab scenario, inspired by wet lab specimen inspection scenarios. In this scenario, two 3-axis gantry robots inspect specimens within a sealed container, as illustrated in Fig. 5(h). The process begins with the robots removing the lid of the container and placing it in a predefined sanitization zone for sterilization. They then use tools attached to their end effectors to inspect the specimen. Once the inspection is complete, the lid is returned to the container to preserve the specimen’s environment. Due to the limited space in the sanitization zone, the planner must determine a valid placement that prevents lid collisions. This placement is constrained by vertex-vertex conflicts, where each vertex represents a potential lid position in the sanitization zone. The results demonstrate that Lazy-DaSH outperforms DaSH, successfully handling more than twice the number of objects requiring multi-stage operations.
@@ V-D Analysis and Discussion
This section analyzes and discusses the experiment results, focusing on five distinct capabilities of Lazy-DaSH using the criteria defined in Section V-A.
@@ V-D1 Overview
Across all tested scenarios, Lazy-DaSH demonstrated significantly improved scalability compared to DaSH, as illustrated in Figures 7 and 8. The scalability analysis in [2] shows that the number of objects leads to linear growth in representation size, while the number of robots results in quadratic growth. This makes DaSH particularly inefficient in environments with many robots, where the exhaustive expansion of the search space imposes substantial computational overhead. This limitation was evident in the Sorting scenarios, which are designed to test scalability with respect to the problem size in terms of both the number of robots and objects.
DaSH also struggles in scenarios that require frequent replanning at both the task and motion levels because it primarily incorporates motion level constraints without adapting the higher level task structure. This limitation is particularly evident in environments with complex constraints such as the Wall scenario, where Lazy DaSH’s optimistic motion validation produced infeasible motions and required constraint feedback to handle invalid robot interactions during the replanning phases. Similarly, the Shelf wall scenario, which is strongly constrained by geometric rules, was not solvable by DaSH. In Lazy DaSH, however, such geometric constraints are identified by the task conflict detection layer, which reviews the unvalidated schedule, detects when a manipulator’s grasp configuration (a vertex) collides with a statically placed object (another vertex), and feeds these constraints back into the task query phase.
DaSH also struggles with non monotonic scenarios because it pre samples object poses without reasoning about their necessity, many of which never contribute to the plan. In contrast, Lazy DaSH expands task space elements only when required. For example, in the Stocking scenario, non monotonicity is identified by the task conflict detection layer, which then triggers selective expansion of the task space. In the Lab scenario, vertex–vertex conflicts (e.g., placing two lids in the same spot) are similarly identified and resolved by resampling to ensure valid placements.
Thus, the hierarchical query framework, combined with a targeted constraint feedback mechanism, enables focused refinements of task and motion representations, thereby improving both efficiency and scalability in complex environments.
@@ V-D2 Representation Size
The planning performance of both DaSH and Lazy-DaSH is directly influenced by the size of their representations, which form the foundation for iterative queries. We compare the sizes of $\mathcal{H}_{\mathcal{T}}$ and $\mathcal{H}_{\mathcal{M}}$ in terms of the number of vertices and hyperarcs, as shown in Figure 7.
As shown in the first two columns, the task space representation of Lazy-DaSH is an order of magnitude larger than that of DaSH, due to the inclusion of start and goal object-only task space elements and their associated robot transitions. In contrast, in comparable scenarios such as Sorting, Wall, and Lab, the motion representation in DaSH, requiring costly collision checking, can be up to two orders of magnitude larger than in Lazy-DaSH. This makes it difficult for DaSH to query over the motion representation using its combined task and motion query approach. Lazy-DaSH, on the other hand, benefits from its constraint management system, which selectively expands only the necessary spaces, leading to a more compact and efficient representation. This difference is particularly evident in the rapid growth of motion hyperarcs in DaSH compared to Lazy-DaSH, especially in the Sorting and Wall scenarios, which involve four and five robots, respectively. Task space expansion is also observed in the Stocking scenarios, where non-monotonicity triggers growth in the representation; in these cases, the number of vertices and hyperarcs varies as object poses are resampled to avoid collisions.
The effectiveness of this targeted motion representation expansion and lazy motion validation is further evaluated in terms of planning time in the next section, with their impact on overall planning performance analyzed in the subsequent discussion.
@@ V-D3 Planning Time
Fig. 8 presents the total planning time along with a detailed breakdown. The first three columns represent the key components contributing to the total planning time: task and motion representation construction, task and motion query, and conflict resolution. The fourth column shows the total planning time, which is the sum of these three components. Task conflict detection is included in the total runtime, but it is only on the order of milliseconds, since it relies on simple per-configuration collision checks that are far less costly than motion queries. To facilitate direct comparisons across different planning times, the y-axis for each scenario is kept consistent across all columns.
@@
The representation construction time for both DaSH and Lazy-DaSH are separately shown in task and motion, showing that the motion representation accounting for the largest portion due to the computational cost of collision checking. In DaSH, this process involves constructing a fully validated roadmap upfront. In contrast, Lazy-DaSH adopts a lazy construction approach, initially generating an edge-invalidated roadmap and incrementally refining it in response to failures encountered during motion queries or conflict resolution. In addition to the reduced representation size in Lazy-DaSH, this incremental approach significantly decreases construction time, achieving up to two orders of magnitude reduction in scenarios where both methods are comparable.
@@
The query times for DaSH and Lazy-DaSH differ in definition. In DaSH, it represents the combined task and motion planning query time, whereas in Lazy-DaSH, it is differentiated with task and motion queries. In both methods, queries may be iteratively triggered by failures in planning or conflict resolution, and the cumulative time for these iterations is reflected in the total query time. A key observation is that as the number of objects increases, the query time in DaSH grows at a much steeper rate compared to Lazy-DaSH. This is due to DaSH’s representation size becoming intractable, leading to an exponential increase in search space. In contrast, Lazy-DaSH maintains a more scalable query process, significantly reducing query time. The reduction reaches up to three orders of magnitude in scenarios where DaSH is able to find a solution.
@@
Motion conflict resolution, involving iterative detection and resolution of path conflicts, takes a noticeable share of planning time as the number of objects grows. As the length of the motion plan increases, typically due to an increased number of objects, the likelihood of recomputation and revalidation increases, becoming a primary source of computation in the overall planning process. Although this suggests that the conflict resolution layer could become a bottleneck in the planning process as the problem size increases, it is worth noting that this layer can be replaced with off the shelf schedulers or their variants to improve scalability.
@@ VI Conclusion
In this paper, we introduce Lazy-Decomposable State Space Hypergraph (Lazy-DaSH), a novel hypergraph-based approach to multi-robot object rearrangement that extends DaSH. Lazy-DaSH separates task and motion planning into distinct layers connected through a constraint feedback mechanism and employs a lazy motion evaluation strategy, validating only the motions relevant to the candidate task plan to minimize unnecessary computation. The constraint feedback mechanism manages infeasible task orders and motions identified by the dedicated conflict detection layer, enabling dynamic updates and refinements of both task and motion planning representations. This design allows the planner to maintain a concise representation, support efficient representation-driven queries, and incorporate the constraints critical to task completion. Experimental results across five scenarios demonstrate that Lazy-DaSH significantly improves scalability and planning speed in highly constrained environments compared to DaSH. Additionally, one of the experiments is validated through hardware implementation:https://youtu.be/3eHOzTikcXc. As future work, adapting online execution into the Lazy-DaSH framework to enhance robustness in dynamic environments represents a promising research direction.